# An $\ell^{1}$-Plug-and-Play Approach for Magnetic Particle Imaging Using a Zero Shot Denoiser with Validation on the 3D Open MPI Dataset

### v1

@@ Results
In Tab. 4 we display the PSNR and SSIM values of the reconstructions . With decreased levels of preprocessing the results are competitive with the baseline methods (cf. Tab. 2) and comparable with the reconstruction obtained with the $\ell^{1}$-PnP in Experiment 1.
(a) Shape whitened, rSVD K=2000.
(b) Shape not whitened, rSVD K=2000.
(c) Shape not whitened, rSVD K=6859.
Fig. 3: Reconstructions of the shape phantom using $\ell^{1}$-PnP with increasingly lower levels of preprocesssing.
In Fig. 3 a selected slice of the reconstructed shape phantom with different levels of preprocessing are displayed. We observe that the proposed $\ell^{1}$-PnP method is rather stable with respect to different preprocessing choices. Further, the results suggest the possibility of employing the $\ell^{1}$-PnP with less preprocessed data and our parameter selection strategy. In the next experiment we go a step further, not relying on preprocessing nor on the SVD.
3.5 Experiment 3: Reconstruction Without Preprocessing Based on Our Parameter Selection Strategy
In our third experiment we study the proposed $\ell^{1}$-PnP algorithm on not preprocessed data of the OpenMPI data set. In contrast to Experiment 2, we do not rely on any kind of SVD, neither for preprocessing, parameter selection, nor in the algorithmic part (Tikhonov type problem (9)). Employing methods not based on an SVD are important for future application scenarios since the computation of an SVD will become increasingly cumbersome with increasing problem size.
The dimension of the system matrix in this scenario is $151230\times 6859$. In particular, without the SVD we cannot solve the Euler-Lagrange equation (9) directly, so we use the Conjugated Gradient algorithm with a tolerance of $10^{-12}$ and a number of maximum iterations of $10000$
@@ 3.5 Experiment 3: Reconstruction Without Preprocessing Based on Our Parameter Selection Strategy
We choose the parameters according to Sec 2.4. For roughly estimating the smallest singular values $\sigma_{\mathrm{min}}$ (without having a SVD) we use power iteration (cf. Section S5 in the supplementary material) and take into account the order of magnitude, i.e., we approximate $\sigma_{\mathrm{min}}$ by $10^{\lfloor\log_{10}\tau\rfloor}$ where $\tau$ denotes the calculated value; concretely, this results in a value of the estimate for $\sigma_{\mathrm{min}}$ as $10^{4}$.
(a) Shape 10-th xz-slice.
(b) Resolution 10-th xy-slice.
(c) Concentration 13-th xy-slice.
Fig. 4: Selected slices of the reconstructions obtained using the proposed $\ell^{1}$-PnP method with our parameter selection/stopping criterion strategy and no preprocessing of the data.
Shape Ph. Resolution Ph. Concentration Ph.
PSNR 30.63 31.99 37.32
SSIM 0.951 0.749 0.570
stop it. 10 9 12
Table 5: Evaluation of the reconstructions using our parameter selection/stopping strategy and no preprocessing.
@@ Results
Selected slices for each phantom in the Open MPI Dataset are shown in Fig. 4 and the quality evaluation with PSNR and SSIM is reported in Tab. 5. We observe that the result obtained in this experiments without preprocessing are comparable with the results obtained with the preprocessing and the knowledge of the SVD in Experiment 1 (cf. Tab. 2). The reconstruction time takes about 1 minute, 30 seconds of which are taken to move the full matrix $A$ to the GPU and to form the matrix $A^{*}A$. Finally, the reconstruction with the full matrix with the PnP approach here presented lasts about 30 seconds including the computation of the quantity $A^{*}f$.
@@ 4.1 Discussion
We start out to discuss relations to the paper [4] which is to our knowledge the work closest to ours in the literature. The authors of [4] were the first to use a PnP method for MPI reconstruction. They train their denoiser on “MPI friendly” MRI data using data of images of vessels, more precisely, they derived their training data from time-of-flight cerebral magnetic resonance angiograms (MRA) to closely mimic the vascular structures targeted in their MPI scans. In contrast to [4], we use a different approach for denoising. We employ a larger size DNN which has been trained on a larger size data set consisting of natural images. The employed denoiser is not learned or transfer learned on MPI or similar MRI data. Our results underpin that larger architectures with larger sets of training data can generalize well to situations where the data distribution is not exactly as in the training data, in particular for conservative tasks such as denoising. This is particular appealing due to the very limited amount of (publicly) available MPI data. The authors of [4] provide a qualitative and quantitative evaluation on simulated data, and show the benefit of their PnP approach by qualitative evaluation on the resolution phantom of the Open MPI dataset. In contrast we here provided a quantitative and qualitative evaluation of our $\ell^{1}$-PnP approach on the full 3D Open MPI data set. Moreover, we use a different iterative scheme, namely the HQS, as a basis contrasting the ADMM type approach in [4]. Using HQS is appealing due to its simplicity. Finally, we provide strategies for parameter selection within our algorithm.
Secondly, we discuss the relation with the DIP approach. The DIP approach stands out since, to our knowledge, it is the only machine learning approach for which a qualitative and quantitative evalution has been carried out on the Open MPI dataset [14]. The DIP approach is an unsupervised approach acting on a single image only. This is on the one hand appealing since other training data cannot influence the reconstruction and on the other hand no training data is needed which is particularly appealing in the context of MPI. Although not an unsupervised approach, the proposed method shares the property that we need no training data (due to using the denoiser in a zero shot fashion.) However, the training data of the denoiser may influence the reconstruction where halucinating is less prominent in denoising. Moreover, the DIP approach suffers from high computational times whereas the $\ell^{1}$-PnP can converge in a few seconds for preprocessed data and in about 30 seconds for not preprocessed data.
@@ 4.2 Conclusion
In this work, we have developed a PnP approach with $\ell^{1}$-prior for MPI reconstruction using a generic zero-shot denoiser. In particular, we employed a larger size DNN, namely the DRUnet architecture of [70] which has been trained on a larger size data set consisting of natural images. In order to apply the image denoiser on a 3D volume, we employ slicing along the axes and averaging. We have quantitatively and qualitatively evaluated our approach on the full 3D Open MPI data set [39], which presently represents the only publicly available real MPI data set. To that end, we have employed suitable preprocessing, and we have compared our scheme quantitatively and qualitatively with the DIP approach of [14] as well as the Tikhonov regularization. Moreover, we have shown that the $\ell^{1}$-PnP approach proposed works with our parameter selection strategy and is able to perform fast reconstructions with different levels of preprocessing. In particular, we show that comparable results can be obtained without preprocessing and in particular without relying on a SVD.
@@ Acknowledgment
This research was supported by the Hessian Ministry of Higher Education, Research, Science and the Arts within the Framework of the “Programm zum Aufbau eines akademischen Mittelbaus an hessischen Hochschulen". We would like to thank Tobias Kluth for valuable discussions and support concerning the prepocessing of MPI data and on the DIP method.
@@
S1 Description of the Background Scans
A delta probe is positioned at each grid cell that we can label with a triple $(x,y,z)$ for $x,y,z\in\{1,\cdots,19\}$. The scanning procedure proceeds in inverse lexicographic order of the tuples $(x,y,z)$, i.e., for any given $y$ and $z$, the scans are performed along the $x$-axis with $x$ from 1 to 19, then the $y$ variable is increased by 1. When all scans along $x$ are performed for each value $y$, the $z$ label is updated by 1. The background data - a scan without the delta probe - is collected at the beginning of the calibration process and each time the $y$ variable of the cell position is updated. Equivalently, the delta probe is removed and an empty scan is performed each time we complete the $19$ measurements along the $x$-axis for each $y$ and $z$. This many background scans are collected because the background signal is assumed to be dynamic. With this procedure, each scan of the delta probe $\delta_{(x,y,z)}$ is performed after an empty scan $b_{(y,z)}^{-}$ and before and empty scan $b_{(y,z)}^{+}$. We can correct each scan along the $x$-axis with the background correction obtained from the convex combination of $b_{(y,z)}^{-}$ and $b_{(y,z)}^{+}$, i.e., if $s_{(x,y,z)}$ is the signal collected scanning the delta probe in the cell $(x,y,z)$, and $\omega_{x}=\frac{x}{19}$ for $x=1,\dots,19$, then the background corrected signal $s_{(x,y,z)}^{\mathrm{corr.}}$ is:
$s_{(x,y,z)}^{\mathrm{corr.}}=s_{(x,y,z)}-\left[\omega_{x}b_{(y,z)}^{-}+(1-\omega_{x})b_{(y,z)}^{+}\right].$ (18)
S2 Specific Implementation of the SSIM
Given two images $f,g$, the SSIM is defined as
$\mathrm{SSIM}(f,g)=\left[l(f,g)\right]^{\alpha}\cdot\left[c(f,g)\right]^{\beta}\cdot\left[s(f,g)\right]^{\gamma},$
where
$\begin{split}l(f,g)=\frac{2\mu_{f}\mu_{g}+C_{1}}{\mu_{x}^{2}+\mu_{y}^{2}+C_{1}},\\ c(f,g)=\frac{2\sigma_{f}\sigma_{g}+C_{2}}{\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2}},\\ s(f,g)=\frac{\sigma_{fg}+C_{3}}{\sigma_{f}\sigma_{g}+C_{3}},\end{split}$ (19)
are the luminance (l), contrast (c) and structure (s)  [64] and $\mu_{f}$ is the mean value, $\sigma_{f}$ the standard deviation of $f$; $\mu_{g}$ and $\sigma_{g}$ are analogously defined for $g$ and $\sigma_{fg}$ the covariance of $f$ and $g$. Following the implementation of [14], we set $\alpha=\beta=\gamma=1$ and $C_{1}=(0.01\cdot R)^{2}$, $C_{2}=(0.03\cdot R)^{2}$ and $C_{3}=0.5\cdot C_{2}$ where $R=100$, i.e., the concentration of the Delta Sample used during the calibration process.
S3 Description of the DIP
The minimization of the functional in Eq. (17) is performed with $p=1$ as suggested, with the random input $z$ having entries uniformly distributed in $[0,0.7]$ and the same shape of the output $(1,19,19,19)$; the regularizing architecture of $\varphi_{\theta}$ is an autoencoder obtained by not using skip connections in a U-Net with encoder steps that down-sample by a factor of 2 with 64, 128 and 256 channels respectively (the decoder is symmetric). The training of $\varphi_{\theta}$ has been performed minimizing the functional in Eq. (17) for 20000 iterations with Adam [30] with different learning rates $\alpha_{i}=10^{-i}$, for $i=3,4,5$ and momenta setting $\beta=(0.9,0.999)$. The selection strategy of the DIP reconstruction is the one used in [14]: we extract reconstructions after a number of iterations $s\in\{1,2,\dots,10,12,\dots,30,35,\dots,50,60,\dots,150,175,\dots,500,\\ 600,\dots,2000,2500,\dots,5000,6000,\dots 20000\}$. For each iteration $s$ we have computed the $\mathrm{PSNR}_{\mathrm{max}}$ and selected the iteration $s_{\mathrm{opt.}}$ that maximizes its value.
S4 Bounds on the $\ell^{1}$ Parameter $\alpha$
In each iteration, a soft-thresholding (proximal mapping of the $\ell^{1}$ norm) of parameter $\alpha/\mu_{k}$ is performed (cf. Eq. (11)). Let $u$ be the ground truth particle distribution and $u_{\mathrm{rec}}$ its reconstruction using the system matrix $A$. The matrix $A$ is obtained by scanning a delta distribution with a known concentration of $c_{\delta}$ mmol. It follows that if $c_{u}$ is the maximum concentration level of the ground truth $u$, the $\ell^{\infty}$-norm of the reconstruction is bounded from above $\lVert u_{\mathrm{rec}}\rVert_{\infty}\leq c_{u}/c_{\delta}=:c_{\mathrm{rec}}$. Therefore, to avoid that the soft-thresholding totally compresses the reconstruction, it is necessary that $\alpha/\mu_{k}<c_{\mathrm{rec}}$, i.e., that $\alpha<c_{\mathrm{rec}}\cdot\mu_{0}$. In particular, if some knowledge about the concentration injected and the concentration of the delta distribution is available, it is possible to set a rough bound on the parameter $\alpha$, for example $\alpha<5\%\cdot c_{\mathrm{rec}}\cdot\mu_{0}$ if we want to allow compression of at most $5\%$ of the reconstructed concentration level. In the experiments in Section 3, because in the Open MPI Dataset $c_{\delta}=100$mmol and we know that for each phantom $c_{u}\leq 100$mmol, we can deduce that $c_{\mathrm{rec}}\leq 1$ and that $\alpha\leq\mu_{0}$. In particular, we choose to compress below $0.5\%$ of the concentration level and set for all experiments in Section 3 the parameter $\alpha=\frac{5}{1000}\mu_{0}$.
S5 Power Iteration Method for $\sigma_{\mathrm{min}}$
We consider the matrix $A^{*}A$ and use the power iteration to approximate $\sigma_{\mathrm{max}}^{2}$, then we apply the power iteration method once more to the matrix $A^{*}A-\sigma_{\mathrm{max}}^{2}\mathrm{Id_{N}}$ to approximate its eigenvalue with greatest norm, i.e., $\tilde{\sigma}_{\mathrm{max}}^{2}=\sigma_{\mathrm{min}}^{2}-\sigma_{\mathrm{max}}^{2}$ and can retrieve an approximation of the smallest singular value $\sigma_{\mathrm{min}}^{2}=\tilde{\sigma}_{\mathrm{max}}^{2}+\sigma_{\mathrm{max}}^{2}$.
S6 Results of Experiment 1

### v2

@@ Results
In Table 5 we display the PSNR and SSIM values of the reconstructions. With decreased levels of preprocessing the results are competitive with the baseline methods (cf. Table 2) and comparable with the reconstruction obtained with the ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP in Experiment 1.
(a) ZeroShot-PnP whitened, rSVD K=2000.
(b) ZeroShot-PnP not whitened, rSVD K=2000.
(c) ZeroShot-PnP not whitened, rSVD K=6859.
(d) ZeroShot-$\ell^{1}$-PnP whitened, rSVD K=2000.
(e) ZeroShot-$\ell^{1}$-PnP not whitened, rSVD K=2000.
(f) ZeroShot-$\ell^{1}$-PnP not whitened, rSVD K=6859.
Figure 5: Reconstructions of the shape phantom using the ZeroShot-PnP and ZeroShot-$\ell^{1}$-PnP methods with increasingly lower levels of preprocessing. To properly compare the results we display the 10-th xz-slice as in Figure 1.
In Figure 5 a selected slice of the reconstructed shape phantom with different levels of preprocessing are displayed. We observe that the proposed ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP methods are rather stable with respect to different preprocessing choices. Further, the results suggest the possibility of employing the proposed methods with less preprocessed data. In the next experiment we go a step further, not relying on preprocessing nor on the SVD.
3.6 Experiment 4: Reconstruction With neither Preprocessing nor SVD
In our fourth and last experiment we study the proposed ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP algorithms on the OpenMPI data set with the minimum amount of preprocessing possible. In contrast to Experiment 2, we do not rely on any kind of SVD, neither for preprocessing, parameter selection, nor in the algorithmic part (Tikhonov type problem (9)). Employing methods not based on an SVD is important for future application scenarios since the computation of an SVD will become increasingly cumbersome with increasing problem size. The dimension of the system matrix in this scenario is $151230\times 6859$. In particular, without the SVD we cannot solve the Euler-Lagrange equation (9) directly, so we use the Conjugated Gradient algorithm with a tolerance of $10^{-12}$ and a number of maximum iterations of $10000$
@@ 3.6 Experiment 4: Reconstruction With neither Preprocessing nor SVD
We choose the parameters according to Section 2.4 and validate the number of iterations $n_{\mathrm{it}}$ and the parameter $\mu_{0}$. For the ZeroShot-PnP method the validated parameters are $n_{\mathrm{it}}=9$ and $\mu_{0}=10^{9}$ whereas for the ZeroShot-$\ell^{1}$-PnP $n_{\mathrm{it}}=48$ and $\mu_{0}=5\cdot 10^{10}$.
(a) ZeroShot-$\ell^{1}$-PnP Shape 10-th xz-slice.
(b) ZeroShot-$\ell^{1}$-PnP Resolution 10-th xy-slice.
(c) ZeroShot-$\ell^{1}$-PnP Concentration 13-th xy-slice.
(d) ZeroShot-PnP Shape 10-th xz-slice.
(e) ZeroShot-PnP Resolution 10-th xy-slice.
(f) ZeroShot-PnP Concentration 13-th xy-slice.
Figure 6: Selected slices of the reconstructions obtained using the proposed ZeroShot-$\ell^{1}$-PnP method (upper row) and with the ZeroShot-PnP method (lower row) with the validated parameters and no preprocessing of the data but the exclusion of the frequencies up to $80$ kHz.
Method Metric Shape Ph. Resolution Ph. Concentration Ph.
ZeroShot-PnP $\mathrm{PSNR}_{\mathrm{max}}$ 31.16 32.06 36.97
$\mathrm{SSIM}_{\mathrm{max}}$ 0.953 0.750 0.551
ZeroShot-$\ell^{1}$-PnP $\mathrm{PSNR}_{\mathrm{max}}$ 29.26 31.27 37.36
$\mathrm{SSIM}_{\mathrm{max}}$ 0.942 0.648 0.576
Table 6: Evaluation of the reconstructions using the validated parameters on the Open MPI dataset without preprocessing of the data.
@@ Results
Selected slices for each phantom in the Open MPI Dataset are shown in Figure 6 and the quality evaluation with $\mathrm{PSNR}_{\mathrm{max}}$ and $\mathrm{SSIM}_{\mathrm{max}}$ is reported in Table 6. We observe that the result obtained in this experiments without preprocessing are comparable with the results obtained with the preprocessing and the knowledge of the SVD in Experiment 1 (cf. Table 2). The reconstruction time takes about 90 seconds, 30 seconds of which are taken to move the full matrix $A$ to the GPU and to form the matrix $A^{*}A$. Finally, the reconstruction with the full matrix with the ZeroShot-PnP (with and without the $\ell^{1}$-prior) approach here presented lasts about 1 minute including the computation of the quantity $A^{*}f$.
@@ 4.1 Discussion
We start out to discuss relations to the paper [4] which is to our knowledge the work closest to ours in the literature. The authors of [4] were the first to use a PnP method for MPI reconstruction. They train their denoiser on “MPI-friendly” MRI data using data of images of vessels, more precisely, they derived their training data from time-of-flight cerebral magnetic resonance angiograms (MRA) to closely mimic the vascular structures targeted in their MPI scans. In contrast to [4], we use a different approach for denoising. We employ a larger size neural network which has been trained on a larger size data set consisting of natural images. The employed denoiser is not learned or transfer learned on MPI or similar MRI data. Our results underpin that larger architectures with larger sets of training data can generalize well to situations where the data distribution is not exactly as in the training data, in particular for conservative tasks such as denoising. This is particularly appealing due to the very limited amount of (publicly) available MPI data. The authors of [4] provide a qualitative and quantitative evaluation on simulated data, and show the benefit of their PnP approach by qualitative evaluation on the resolution phantom of the Open MPI dataset. In contrast we provide a quantitative and qualitative evaluation of our ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP approach on the full 3D Open MPI dataset. Moreover, we use a different iterative scheme, namely the HQS, as a basis contrasting the ADMM type approach in [4]. Using HQS is appealing due to its simplicity and the ensuing relationship between the parameters and the Gaussian noise levels. Finally, we validate the parameters on a hybrid dataset. In addition, we discuss the relation with the DIP approach and the previous PP-MPI algorithm with trained denoiser [4]. The DIP approach stands out since, to our knowledge, it is the only machine learning approach for which a qualitative and quantitative evalution has been carried out on the Open MPI dataset [14]. The DIP approach is an unsupervised approach acting on a single image only. This is on the one hand appealing since other training data cannot influence the reconstruction and on the other hand no training data is needed which is particularly appealing in the context of MPI. Although not an unsupervised approach, the proposed method shares the property that we need no training data (due to using the denoiser in a zero shot fashion). However, the training data of the denoiser may influence the reconstruction. Moreover, the DIP approach suffers from high computational times whereas the ZeroShot-$\ell^{1}$-PnP can converge in a few seconds for preprocessed data and in about one minute for not preprocessed data. Finally, the benefit of using the zero-shot denoiser in the context of MPI is not only suggested by the quality of both the qualitative and quantitative results but also by its flexibility; in fact, in the previous PP-MPI approach the denoiser has been trained on MPI-friendly images of size $32\times 32$ pixels and with training data ranging between 0 and $1.5$. This means that, if the dimensionality of the MPI system-matrix inversion problem increases, it is reasonable to expect that another ad hoc denoiser shall be trained on a new dataset in order to be able to denoise more complicated images. The usage of a zero-shot denoiser pre-trained on a wide range of natural images might enable the employment of the suggested ZeroShot-PnP algorithm with little additional changes.
@@ 4.2 Conclusion
In this work, we have developed a PnP approach with $\ell^{1}$-prior for MPI reconstruction using a generic zero-shot denoiser. In particular, we employed a larger size DNN, namely the DRUnet architecture of [70] which has been trained on a larger size data set consisting of natural images. In order to apply the image denoiser on volumetric data, we employ slicing along the axes and averaging. We have quantitatively and qualitatively evaluated our approach on the full 3D Open MPI data set [39], which presently represents the only publicly available real MPI data set. To that end, we have employed suitable preprocessing, and we have compared our scheme quantitatively and qualitatively with the DIP approach of [14] as well as the Tikhonov regularization and the previous Plug-and-Play algorithm with trained denoiser [4]. Moreover, we have shown that the ZeroShot-$\ell^{1}$-PnP approach proposed works with parameters validated on a hybrid dataset and is able to perform fast reconstructions with different levels of preprocessing. In particular, we show that comparable results can be obtained without preprocessing and in particular without relying on SVD of the matrix.
@@ Acknowledgment
This research was supported by the Hessian Ministry of Higher Education, Research, Science and the Arts within the Framework of the “Programm zum Aufbau eines akademischen Mittelbaus an hessischen Hochschulen" and by the German Science Fonds DFG under grant INST 168/4-1. We would like to thank Tobias Kluth for valuable discussions and support concerning the preprocessing of MPI data and the DIP method.